# Mechanical quadrature squeezing beyond the 3dB limit via quantum learning control

### v1

@@ 3 Generation of mechanical quadrature squeezing
A positive squeezing degree, $\mathcal{S}_{b}>0$, implies that the quadrature operator of the mechanical mode is squeezed along the angle $\theta$. For studying the rotating-quadrature squeezing, we need to investigate the dependence of the squeezing on the angle $\theta$. Concretely, we use the quantum-learning-control method to obtain the waveforms of the driving amplitude and phase corresponding to $\mathcal{S}_{b}=1$ at different values of $\theta$. The results shown in Fig. 1(a) indicate that the smallest (largest) driving amplitude appears at the squeezing angle $90^{\circ}$ ($0^{\circ}$) related to $\mathcal{S}_{b}=1$. Similarly, the minimal and maximal driving phases appear at $65^{\circ}$ and $165^{\circ}$, respectively, and the angle has little effect on the phase [as shown in Fig. 1(b)]. Therefore, we take the squeezing angle $\theta=90^{\circ}$ (corresponding to minimal driving amplitude) in our following discussions.
We display in Fig. 2(a) the loss function $\Delta X_{b}^{2}(\pi/2,T)$ as a function of the iteration number to verify the efficiency of the quantum-learning-control method. Here, we can see that the loss function $\Delta X_{b}^{2}(\pi/2,T)$ decreases gradually as the iteration number increases. The transient mechanical squeezing degrees 1 dB and 3.2 dB are obtained at the iteration numbers 1185 and 4000, respectively. The corresponding dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ are shown in Figs. 2(b) and 2(c). It shows that there are no mechanical squeezing for a long duration of time firstly, and then oscillatingly increase to the target value in a short period of time. In particular, the squeezing result $\mathcal{S}_{b}=3.2$ breaks the 3 dB steady-state limit [Fig. 2(c)], and greater squeezings can be realized as the iteration number increases. Note that we consider the transient squeezing here rather than steady-state squeezing. Therefore, our results are not conflicting with the 3 dB steady-state limit, which is determined by the dynamic stability.
To confirm the mechanical squeezing in phase space, we introduce the Wigner function of the mechanical mode, defined as [62]
@@
where $\mathbf{D}=[D_{\mathrm{R}},D_{\mathrm{I}}]^{\mathrm{T}}$ represents the two-dimensional vector, and $\mathbf{V}_{b}$ is the covariance matrix for the mechanical mode. The Wigner functions corresponding to 1 dB and 3.2 dB squeezing are shown in Figs. 2(d) and 2(e). We can see that the squeezing appears along the $\pi/2$-axis (corresponding to the squeezing angle $\theta=\pi/2$), and the larger the squeezing degree, the stronger the quadrature squeezing.
In Figs. 2(f) and 2(g), we show the time-dependent driving amplitudes, which are required to achieve the squeezing degrees $\mathcal{S}_{b}=1$ dB and 3.2 dB, respectively [^1] We checked the correctness of the pulsed fields by back-substituting the pulsed field data obtained after the last iteration into the system evolution. We verified the results by either solving the cavariance matrix $\mathbf{V}$ or solving the quantum master equation (4) under the replacement of $H_{\mathrm{dis}}(t)\rightarrow H_{\mathrm{lin}}(t)$. We find that the dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ are consistent with the results obtained in this paper.. We can see that a larger driving amplitude is required to realize a stronger squeezing. In addition, for a larger mechanical squeezing degree, the driving phase oscillation becomes more intense, as shown in Figs. 2(h) and 2(i). We also investigate the dependence of the mean phonon number $\langle b^{\dagger}b\rangle$ (in the displaced representation, associated with the quantum fluctuation) as a function of the scaled time $\omega_{\mathrm{m}}t$ in Figs. 2(j) and 2(k). We observe that the $\langle b^{\dagger}b\rangle$ decreases from one hundred to less than one for both $\mathcal{S}_{b}=1$ dB and 3.2 dB. Note that these results can be well understood by examining the statistical properties for the squeezed vacuum state. For a single-mode squeezed vacuum state, the mean phonon number is $\sinh^{2}r$, which is a function of the squeezing amplitude $r$. Further, the covariance of the mechanical rotating-quadrature operator in the squeezing-vacuum state is $(2\sinh^{2}r+1-2\sinh r\cosh r\cos\varphi)/2$, ($\varphi$ is the difference between the rotation angle and the squeezing angle). Considering $\varphi=0$, then we can numerically verify that the squeezing amplitude corresponding to the squeezing degree $\mathcal{S}_{b}=3$ is $r=0.347$, and the mean phonon number is 0.125. This indicates that the mechanical squeezing exists even if the mechanical resonator is cooled near the ground state.
@@
Ultrafast generation of mechanical squeezing is a desired task from the viewpoint of transient evolution. Below, we investigate the generation of mechanical squeezing within one mechanical oscillation period. Figure 4(a) shows the loss function $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ versus the iteration number related to $\mathcal{S}_{b}=1$. Here we can see that the $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ gradually decreases from the initial value 319 to 0.397 as the iteration number increases, verifying the feasibility of the gradient descent algorithm. The dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ after the last iteration ($m=6180$) is shown in Fig. 4(b). We see that the squeezing occurs in the last extremely short duration due to the harsh time condition, but it is still continuous, as shown in the inset of Fig. 4(b). The results in Figs. 4(c) and 4(d) indicate that a large driving amplitude is required to achieve $\mathcal{S}_{b}=1$ within one mechanical oscillation period and the phase has no distinct signature. We also exhibit in Fig. 4(e) the mean phonon number $\langle b^{\dagger}b\rangle$ versus the scaled evolution time at this time. Here we see that $\langle b^{\dagger}b\rangle$ will be larger than the initial value due to the larger driving amplitude, and then goes to a cooled state with dozens of mean phonons. This indicates that the creation of mechanical squeezing does not necessarily require the ground-state cooling of the mechanical resonator.
@@
We also investigate the influence of the sideband-resolution parameter on the squeezing generation. In Figs. 5(a-c), we plot the squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to different sideband-resolution parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.1. We can see that the deterioration of the sideband-resolution parameter will reduce the squeezing-evolution peak. Furthermore, we show the driving amplitude and phase as functions of $\omega_{\mathrm{m}}t$ for different sideband-resolution parameters in Figs. 5(d-f) and Figs. 5(g-i). To maintain $\mathcal{S}_{b}=1$, a larger driving amplitude is needed for a larger value of $\kappa/\omega_{\mathrm{m}}$, and the phase oscillation becomes more severe for a larger $\kappa/\omega_{\mathrm{m}}$. The corresponding dynamic evolution of the mean phonon number $\langle b^{\dagger}b\rangle$ is shown in Fig. 5(j). Here we can see that the mechanical resonator is cooled from the initial one hundred phonons to the ground state. The better the sideband-resolution parameter is selected, the deeper cooling of the mechanical resonator.
@@
Figure 5: (a-c) The squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ under different sideband parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.1. The inset in panels (b,c) exhibit the evolution of $\mathcal{S}_{b}$ for the time durations [115.6,120] and [116.2,120], respectively. (d-f) Evolution of driving amplitude and (g-i) phase corresponding to panels (a-c). (j) The dynamic evolution of the mean phonon number under different sideband-resolution parameters in panels (a-c). Other parameters are consistent with those given in Fig. 1.
4 Discussions and conclusion
@@
$\Delta X_{b}^{2}(\theta,t)=\frac{1}{2}+\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta).$ (15)
@@
Since the mechanical squeezing occurs when $\Delta X_{b}^{2}(\theta,t)<1/2$, our goal is to realize $\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0$, which is feasible because the second-order moment $\langle b^{\dagger}b^{\dagger}\rangle$ could be a complex number. Meanwhile, the value of $\langle b^{\dagger}b\rangle$ must be a positive number, and then the reduction of $\langle b^{\dagger}b\rangle$ is beneficial to the squeezing generation, which explains why the mechanical resonator is cooled [Figs. 3(c), 3(f), 4(e), and 5(j)]. In addition, under the rotating-wave approximation, we have two sets of equations governing all these second-order moments. The first set includes $\langle a^{\dagger}a\rangle$, $\langle b^{\dagger}b\rangle$, $\langle a^{\dagger}b\rangle$, and $\langle ab^{\dagger}\rangle$, while the second set includes $\langle a^{\dagger}a^{\dagger}\rangle$, $\langle aa\rangle$, $\langle b^{\dagger}b^{\dagger}\rangle$, $\langle bb\rangle$, $\langle a^{\dagger}b^{\dagger}\rangle$, and $\langle ab\rangle$. Initially, with the condition $\mathbf{X}(0)=[0,\bar{n}_{\mathrm{m}},0,0,0,0,0,0,0,0]^{\mathrm{T}}$, the value of $\langle b^{\dagger}b^{\dagger}\rangle$ remains constantly zero, making mechanical squeezing unattainable regardless of how we optimize the driving field. In the presence of the counter-rotating terms, the equations of motion for the ten second-order moments are coupled to each other. Consequently, $\langle b^{\dagger}b^{\dagger}\rangle$ is no longer zero. In this case, we can obtain $\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0$ via optimizing the pulsed driving. These analyses indicate that the counter-rotating terms play a crucial role in the generation of mechanical quadrature squeezing.
@@
We now present some discussions on the experimental feasibility of this scheme. The system under considerations is a general optomechanical system, which has been implemented in various optomechanical platforms, such as optical microresonators [63, 64, 65], electromechanical systems [66, 67, 68], photonic crystal nanobeams [5, 69, 70], and Fabry-Pérot cavities [71], in which the cavities can be driven by tunable pulsed fields. We considered the linearized optomechanical interaction, which has been widely demonstrated in cavity optomechanical systems. Moreover, we use the experimentally accessible parameters in our numerical simulations. Concretely, the used parameters are $g_{0}/\omega_{\mathrm{m}}=4\times 10^{-5}$, $\kappa/\omega_{\mathrm{m}}=0.2$, and $\gamma/\omega_{\mathrm{m}}=2\times 10^{-6}$, which have been reported in experiments [65]. In addition, the pulsed driving amplitude and phase used to generate mechanical squeezing are moderate in size and continuous and smooth in shape, which confirm the experimental realization of the pulsed driving.
@@
In conclusion, we have presented a scheme for generating mechanical quadrature squeezing in a typical optomechanical system via a quantum-learning-control method. The generated mechanical squeezing can exceed the 3 dB steady-state limit and ultrafast squeezing preparation within one mechanical oscillation period can be realized. The optimal driving amplitude and phase corresponding to these generated squeezings have been presented. Our scheme will pave the way for exploiting quantum learning control in quantum optics and quantum information science.
Code availability
The source code is available upon reasonable request from the authors.
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary files).
@@
The authors thank Ye-Xiong Zeng and Yue-Hui Zhou for helpful discussions. J.-Q.L. was supported in part by National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, and No. 11935006), the Science and Technology Innovation Program of Hunan Province (Grant No. 2021RC4029), and Hunan Provincial Major Science and Technology Program (Grant No. 2023ZJ1010).
Appendix. Derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$ in gradient descent algorithm

### v2

@@ III GENERATION OF MECHANICAL QUADRATURE SQUEEZING
A positive squeezing degree, $\mathcal{S}_{b}>0$, implies that the quadrature operator of the mechanical mode is squeezed along the angle $\theta$. For studying the rotating-quadrature squeezing, we need to investigate the dependence of the squeezing on the angle $\theta$. Concretely, we use the gradient-descent algorithm to obtain the waveforms of the driving amplitude and phase corresponding to $\mathcal{S}_{b}=1$ at different values of $\theta$. The results shown in Fig. 1(a) indicate that the smallest (largest) driving amplitude appears at the squeezing angle $90^{\circ}$ ($0^{\circ}$) related to $\mathcal{S}_{b}=1$. Similarly, the minimal and maximal driving phases appear at $65^{\circ}$ and $165^{\circ}$, respectively, and the angle has little effect on the phase [as shown in Fig. 1(b)]. Therefore, we take the squeezing angle $\theta=90^{\circ}$ (corresponding to minimal driving amplitude) in our following discussions.
We display in Fig. 2(a) the loss function $\Delta X_{b}^{2}(\pi/2,T)$ as a function of the iteration number to verify the efficiency of the gradient-descent algorithm. Here, we can see that the loss function $\Delta X_{b}^{2}(\pi/2,T)$ decreases gradually as the iteration number increases. The transient mechanical squeezing degrees $1$ dB and $3.2$ dB are obtained at the iteration numbers $1185$ and $4000$, respectively. The corresponding dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ are shown in Figs. 2(b) and 2(c). It shows that there are no mechanical squeezing for a long duration of time firstly, and then oscillatingly increase to the target value in a short period of time. In particular, the squeezing degree $\mathcal{S}_{b}=3.2$ breaks the $3$ dB steady-state limit [Fig. 2(c)], and stronger squeezings can be realized as the iteration number increases. Note that we consider the transient squeezing here rather than steady-state squeezing. Therefore, our results are not conflicting with the $3$ dB steady-state squeezing limit, which is determined by the dynamic stability.
To confirm the mechanical squeezing in phase space, we introduce the Wigner function of the mechanical mode, defined as Vogel2006
@@
where $\mathbf{D}=(D_{\mathrm{R}},D_{\mathrm{I}})^{\mathrm{T}}$ represents the two-dimensional vector, and $\mathbf{V}_{b}$ is the covariance matrix for the mechanical mode. The Wigner functions corresponding to 1 dB and 3.2 dB squeezing are shown in Figs. 2(d) and 2(e). We can see that the squeezing appears along the $\pi/2$-axis (corresponding to the squeezing angle $\theta=\pi/2$), and the larger the squeezing degree, the stronger the quadrature squeezing.
In Figs. 2(f) and 2(g), we show the time-dependent driving amplitudes, which are required to achieve the squeezing degrees $\mathcal{S}_{b}=1$ dB and 3.2 dB, respectively check . We can see that a larger driving amplitude is required to realize a stronger squeezing. In addition, for a larger mechanical squeezing degree, the driving phase oscillation becomes more intense, as shown in Figs. 2(h) and 2(i). We also investigate the dependence of the mean phonon number $\langle b^{\dagger}b\rangle$ (in the displaced representation, associated with the quantum fluctuation) as a function of the scaled time $\omega_{\mathrm{m}}t$ in Figs. 2(j) and 2(k). We observe that $\langle b^{\dagger}b\rangle$ decreases from one hundred to less than one for both $\mathcal{S}_{b}=1$ dB and 3.2 dB. Note that these results can be well understood by examining the statistical properties for the squeezed vacuum state. For a single-mode squeezed vacuum state, the mean phonon number is $\sinh^{2}r$, which is a function of the squeezing amplitude $r$. Further, the covariance of the mechanical rotating-quadrature operator in the squeezing-vacuum state is $(2\sinh^{2}r+1-2\sinh r\cosh r\cos\varphi)/2$, ($\varphi$ is the difference between the rotation angle and the squeezing angle). Considering $\varphi=0$, then we can numerically verify that the squeezing amplitude corresponding to the squeezing degree $\mathcal{S}_{b}=3$ is $r=0.347$, and the mean phonon number is 0.125. This indicates that the mechanical squeezing exists even if the mechanical resonator is cooled near the ground state.
@@
Ultrafast generation of mechanical squeezing is a desired task from the viewpoint of transient evolution. Below, we investigate the generation of mechanical squeezing within one mechanical oscillation period. Figure 4(a) shows the loss function $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ versus the iteration number related to $\mathcal{S}_{b}=1$. Here we can see that the $\Delta X_{b}^{2}(\pi/2,T=6\omega_{\mathrm{m}}^{-1})$ gradually decreases from the initial value 319 to 0.397 as the iteration number increases, verifying the validity of the gradient-descent algorithm. The dynamic evolution of the squeezing degree $\mathcal{S}_{b}$ after the last iteration ($m=6180$) is shown in Fig. 4(b). We see that the squeezing occurs in the last extremely short duration due to the harsh time condition, but it is still continuous, as shown in the inset of Fig. 4(b). The results in Figs. 4(c) and 4(d) indicate that a large driving amplitude is required to achieve $\mathcal{S}_{b}=1$ within one mechanical oscillation period and the phase has no distinct signature. We also exhibit in Fig. 4(e) the mean phonon number $\langle b^{\dagger}b\rangle$ versus the scaled evolution time at this time. Here we see that $\langle b^{\dagger}b\rangle$ will be larger than the initial value due to the larger driving amplitude, and then goes to a cooled state with dozens of mean phonons. This indicates that the creation of mechanical squeezing does not necessarily require the ground-state cooling of the mechanical resonator.
@@
We also investigate the influence of the sideband-resolution parameter on the squeezing generation. In Figs. 5(a-c), we plot the squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ corresponding to different sideband-resolution parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.5. We show the driving amplitude and phase as functions of $\omega_{\mathrm{m}}t$ for different sideband-resolution parameters in Figs. 5(d-f) and Figs. 5(g-i). To maintain $\mathcal{S}_{b}=1$, a larger driving amplitude is needed for a larger value of $\kappa/\omega_{\mathrm{m}}$, and the phase oscillation becomes more severe for a larger $\kappa/\omega_{\mathrm{m}}$. The corresponding dynamic evolution of the mean phonon number $\langle b^{\dagger}b\rangle$ is shown in Fig. 5(j). Here we can see that the mechanical resonator is cooled from the initial one hundred phonons to a few phonons. The better the sideband-resolution parameter is selected, the deeper cooling of the mechanical resonator. Note that a large $\kappa/\omega_{\mathrm{m}}$ will result in fast photon dissipation within the cavity, thereby affecting the optomechanical interaction and leading to the failure of squeezing generation.
@@
Figure 5: (a-c) The squeezing degree $\mathcal{S}_{b}$ versus the scaled evolution time $\omega_{\mathrm{m}}t$ under different sideband parameters $\kappa/\omega_{\mathrm{m}}=0.5$, 1.0, and 1.5. The inset in panel (b) exhibits the evolution of $\mathcal{S}_{b}$ for the time durations from $115.6$ to $120$. (d-f) Evolution of driving amplitude and (g-i) phase corresponding to panels (a-c). (j) The dynamic evolution of the mean phonon number under different sideband-resolution parameters in panels (a-c). Other parameters are consistent with those given in Fig. 1.
IV DISCUSSIONS AND CONCLUSION
@@
$\Delta X_{b}^{2}(\theta,t)$ $=\frac{1}{2}+\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)$
$\hskip 8.5359pt-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta).$ (15)
@@
Since the mechanical squeezing occurs when $\Delta X_{b}^{2}(\theta,t)<1/2$, our goal is to realize $\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0$, which is feasible because the second-order moment $\langle b^{\dagger}b^{\dagger}\rangle$ could be a complex number. Meanwhile, the value of $\langle b^{\dagger}b\rangle$ must be a positive number, and then the reduction of $\langle b^{\dagger}b\rangle$ is beneficial to the squeezing generation, which explains why the mechanical resonator is cooled [Figs. 2(j), 2(k), 4(e), and 5(j)]. In addition, under the rotating-wave approximation, we have two sets of equations governing all these second-order moments. The first set includes $\langle a^{\dagger}a\rangle$, $\langle b^{\dagger}b\rangle$, $\langle a^{\dagger}b\rangle$, and $\langle ab^{\dagger}\rangle$, while the second set includes $\langle a^{\dagger}a^{\dagger}\rangle$, $\langle aa\rangle$, $\langle b^{\dagger}b^{\dagger}\rangle$, $\langle bb\rangle$, $\langle a^{\dagger}b^{\dagger}\rangle$, and $\langle ab\rangle$. Initially, with the condition $\mathbf{X}(0)=[0,\bar{n}_{\mathrm{m}},0,0,0,0,0,0,0,0]^{\mathrm{T}}$, the value of $\langle b^{\dagger}b^{\dagger}\rangle$ remains constantly zero, making mechanical squeezing unattainable regardless of how we optimize the driving field. In the presence of the counter-rotating terms, the equations of motion for these ten second-order moments are coupled to each other. Consequently, $\langle b^{\dagger}b^{\dagger}\rangle$ is no longer zero. In this case, we can obtain $\langle b^{\dagger}b\rangle+\mathrm{Re}[\langle b^{\dagger}b^{\dagger}\rangle]\cos(2\theta)-2\mathrm{Im}[\langle b^{\dagger}b^{\dagger}\rangle]\sin(2\theta)<0$ via optimizing the pulsed driving. These analyses indicate that the counter-rotating terms play a crucial role in the generation of mechanical quadrature squeezing.
@@
We now present some discussions on the experimental feasibility of this scheme. The system under considerations is a general optomechanical system, which has been implemented in various optomechanical platforms, such as optical microresonators ParkNP2009 ; SchliesserNP2009 ; VerhagenNature2012 , electromechanical systems HertzbergNP2010 ; MasselN2011 ; MasselNC2012 , photonic crystal nanobeams TeufelNature2011 ; Eichenfield2009 ; ChanNT2009 , and Fabry-Pérot cavities HammererNature2009 , in which the cavities can be driven by tunable pulsed fields. We considered the linearized optomechanical interaction, which has been widely demonstrated in cavity optomechanical systems. Moreover, we use the experimentally accessible parameters in our numerical simulations. Concretely, the used parameters are $g_{0}/\omega_{\mathrm{m}}=4\times 10^{-5}$, $\kappa/\omega_{\mathrm{m}}=0.2$, and $\gamma/\omega_{\mathrm{m}}=2\times 10^{-6}$, which have been reported in experiments VerhagenNature2012 . In addition, the pulsed driving amplitude and phase used to generate squeezing are moderate in size and continuous and smooth in shape, which confirm the experimental realization of the pulsed driving.
@@
In conclusion, we have presented a scheme for generating mechanical quadrature squeezing in a typical optomechanical system via gradient-descent algorithm. The generated mechanical squeezing can exceed the 3 dB steady-state limit and ultrafast squeezing preparation within one mechanical oscillation period can be realized. The optimal driving amplitude and phase corresponding to these generated squeezings have been presented. Our scheme will pave the way for exploiting optimal quantum control in quantum optics and quantum information science.
Acknowledgements.
@@
The authors thank Dr. Ye-Xiong Zeng and Dr. Yue-Hui Zhou for helpful discussions. J.-Q.L. was supported in part by National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, and No. 11935006), the Science and Technology Innovation Program of Hunan Province (Grant No. 2021RC4029), and Hunan Provincial Major Science and Technology Program (Grant No. 2023ZJ1010).
Appendix A Derivation of the variation $\delta[\Delta X_{b}^{2}(\theta,T)]/\delta\mathcal{Q}_{m}$